# Explicit Performance Bound of Finite Blocklength Coded MIMO: Time-Domain versus Spatiotemporal Channel Coding

### v2

@@ Appendix B Proof of Theorem 1
According to the derivation procedure of the normal approximation for a quasi-static flat Rayleigh fading MIMO channel in (31), the normal approximation is accurate when
@@ B-A Large Blocklength
$n{{C}^{\text{ST}}}\left(\mathsf{H}\right)-\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\varepsilon\right)=n\left[{{C}^{\text{ST}}}\left(\mathsf{H}\right)-\sqrt{\frac{{{V}^{\text{ST}}}\left(\mathsf{H}\right)}{n}}\Phi^{-1}\left(\varepsilon\right)\right].$ (64)
@@ B-D Large Spatial DoF
$n{{C}_{m+1}^{\text{ST}}}\left(\mathsf{H}\right)-\sqrt{n{{V}_{m+1}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\varepsilon\right)-\left[n{{C}_{m}^{\text{ST}}}\left(\mathsf{H}\right)-\sqrt{n{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\varepsilon\right)\right]$
$=n{\log\left(1+\frac{\rho}{L}{{\lambda}_{m+1}}\right)}+\sqrt{n{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\varepsilon\right)-\sqrt{n\left[{\left(1-\frac{1}{{{\left(1+\frac{\rho}{L}{{\lambda}_{m+1}}\right)}^{2}}}\right)\log^{2}e+{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}\right]}\Phi^{-1}\left(\varepsilon\right)$
@@
$\sqrt{n{\left(1-\frac{1}{{{\left(1+\frac{\rho}{L}{{\lambda}_{m+1}}\right)}^{2}}}\right)\log^{2}e}}>-\sqrt{n{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}+\sqrt{n\left[{\left(1-\frac{1}{{{\left(1+\frac{\rho}{L}{{\lambda}_{m+1}}\right)}^{2}}}\right)\log^{2}e+{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}\right]},$ (73)
then (B-D) must hold. In this case, we should prove that
$\sqrt{n{\left(1-\frac{1}{{{\left(1+\frac{\rho}{L}{{\lambda}_{m+1}}\right)}^{2}}}\right)\log^{2}e}}+\sqrt{n{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}>\sqrt{n\left[{\left(1-\frac{1}{{{\left(1+\frac{\rho}{L}{{\lambda}_{m+1}}\right)}^{2}}}\right)\log^{2}e+{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}\right]}.$ (74)
Taking the square of both sides of (74) simultaneously, we have that
${{\left[1-\frac{1}{{{\left(1+\frac{\rho}{L}{{\lambda}_{m+1}}\right)}^{2}}}\right]\log^{2}e}}+{{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}+2\sqrt{\left[1-\frac{1}{{{\left(1+\frac{\rho}{L}{{\lambda}_{m+1}}\right)}^{2}}}\right]\log^{2}e{{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}}$
$>{{\left[1-\frac{1}{{{\left(1+\frac{\rho}{L}{{\lambda}_{m+1}}\right)}^{2}}}\right]\log^{2}e}}+{{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}$ (75)
holds true constantly. This indicates that as the spatial DoF increases, $n{{C}^{\text{ST}}}\left(\mathsf{H}\right)-\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\varepsilon\right)\gg\log n$, and the normal approximation will become more accurate.

### v3

@@ Appendix B Proof of Theorem 1
According to the derivation procedure of the normal approximation for a quasi-static flat Rayleigh fading MIMO channel in (27), when the higher order terms are small enough to be neglected, the normal approximation that preserves the first two order terms is accurate. Since the decay of the higher order terms is going to be less than $\log n$, we just need to figure out when
@@ B-A Large Blocklength
$n{{C}^{\text{ST}}}\left(\mathsf{H}\right)-\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\varepsilon\right)$
$\qquad=n\left[{{C}^{\text{ST}}}\left(\mathsf{H}\right)-\sqrt{\frac{{{V}^{\text{ST}}}\left(\mathsf{H}\right)}{n}}\Phi^{-1}\left(\varepsilon\right)\right].$ (60)
@@ B-D Large Spatial DoF
$n{{C}_{m+1}^{\text{ST}}}\left(\mathsf{H}\right)-\sqrt{n{{V}_{m+1}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\varepsilon\right)$
$\qquad-\left[n{{C}_{m}^{\text{ST}}}\left(\mathsf{H}\right)-\sqrt{n{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\varepsilon\right)\right]$
$=n{\log\left(1+\frac{\rho}{L}{{\lambda}_{m+1}}\right)}+\sqrt{n{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\varepsilon\right)$
$\quad-\sqrt{n\left[{\left(1-\frac{1}{{{\left(1+\frac{\rho}{L}{{\lambda}_{m+1}}\right)}^{2}}}\right)\log^{2}e+{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}\right]}\Phi^{-1}\left(\varepsilon\right)$
@@
$\sqrt{n{\left(1-\frac{1}{{{\left(1+\frac{\rho}{L}{{\lambda}_{m+1}}\right)}^{2}}}\right)\log^{2}e}}+\sqrt{n{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}$
$\quad>\sqrt{n\left[{\left(1-\frac{1}{{{\left(1+\frac{\rho}{L}{{\lambda}_{m+1}}\right)}^{2}}}\right)\log^{2}e+{{V}_{m}^{\text{ST}}}\left(\mathsf{H}\right)}\right]}.$ (69)
Taking the square of both sides of (B-D) simultaneously, we found it holds true constantly. This indicates that as the spatial DoF increases, $n{{C}^{\text{ST}}}\left(\mathsf{H}\right)-\sqrt{n{{V}^{\text{ST}}}\left(\mathsf{H}\right)}\Phi^{-1}\left(\varepsilon\right)\gg\log n$, and the normal approximation will become more accurate.